A sample of nitrogen occupies a volume of 250 mL at 25 C. What volume will it occupy at 0C?


1
Expert's answer
2022-05-05T03:22:07-0400

V1/T1 = V2/T2

V2 = (V1/T1) * T2

V1 = 250 mL = 0.25 L, T1 = 25°C = 298.15 K, T2 = 0°C = 273.15 K

V2 = (0.25 L / 298.15 K) * 273.15 K = 0,229 L = 229  mL